### v2

# TestDG: Test-time Domain Generalization for Continual Test-time Adaptation

Abstract
This paper studies continual test-time adaptation (CTTA), the task of adapting a model to constantly changing unseen domains in testing while preserving previously learned knowledge. Existing CTTA methods mostly focus on adaptation to the current test domain only, overlooking generalization to arbitrary test domains a model may face in the future. To tackle this limitation, we present a novel online test-time domain generalization framework for CTTA, dubbed TestDG. TestDG aims to learn features invariant to both current and previous test domains on the fly during testing, improving the potential for effective generalization to future domains. To this end, we propose a new model architecture and a test-time adaptation strategy dedicated to learning domain-invariant features, along with a new data structure and optimization algorithm for effectively managing information from previous test domains. TestDG achieved state of the art on four public CTTA benchmarks. Moreover, it showed superior generalization to unseen test domains.
1 Introduction
Deep neural networks have driven significant advances in various vision tasks such as classification [24, 40, 13], object detection [20, 6, 5], and semantic segmentation [51, 69, 9]. Despite these achievements, they often struggle with limited generalization ability to the domain shift between training and test data [18, 25, 60, 8, 53, 55, 54, 36, 4, 37]. Test-time adaptation (TTA) [73, 64, 46, 72, 49, 50, 30, 31] has been developed to mitigate the distribution shift by adapting a pre-trained model to unlabeled test domains during testing. In specific, TTA methods update the model in testing on the fly, by self-training using pseudo-labeled test data [58, 17, 43] or by entropy minimization of the model prediction [49, 64]. Early approaches to TTA assume a single and fixed test domain. In reality, however, this assumption does not hold since a model is often deployed in non-stationary and continually changing environments. In such environments, the accuracy of the model can easily degrade even with TTA since the constant distribution shift in testing causes unreliable pseudo labels [22], which in turn exacerbates error accumulation and catastrophic forgetting.
Figure 1: Empirical demonstration of the effectiveness of TestDG. (a) Accuracy across four CTTA and one domain generalization (DG) benchmarks, where DG performance is evaluated on 5 unseen domains after CTTA over 10 domains. For ACDC, mIoU is used instead. The top-2 prior arts are shown as dots with different shapes for each benchmark. TestDG achieved the best on all benchmarks. (b) Error rate over sequentially changing test domains. The consistently low error rates and notably stable performance of TestDG, especially in moments of domain change, highlight its robustness.
Continual test-time adaptation (CTTA) [67, 58, 17, 43, 34, 70, 42, 71, 77] has been studied to address these realistic TTA scenarios. Most of existing CTTA methods focus only on adapting a model to the test domain at hand. Although these methods lead to notable performance improvements, there remains room for further improvement in that they do not fully account for the model’s generalization to future test domains it encounters later in continuously changing environments.
In this paper, we present an online TEST-time Domain Generalization framework for continual test-time adaptation, dubbed TestDG. TestDG overcomes the aforementioned limitation by learning domain-invariant representations that well generalize to unseen domains on the fly during testing. This approach improves robustness of a model to arbitrary test domains that it may encounter in the future, as demonstrated in Figure 1(a). Furthermore, as shown in Figure 1(b), learned domain-generalizable features enable robust adaptation even under abrupt domain shifts. A naïve approach to implementing domain-invariant learning is to align features from different test domains. However, since the features contain their own semantic contents as well as domain-specific information, this alignment without consideration of semantics corrupts the features and consequently degrades performance.
Hence, we introduce domain information extractor that takes the encoder features as input and extracts only domain-specific information in the form of embedding vectors; we call such vectors domain embeddings. The encoder is, in turn, trained to make domain embeddings from different domains indistinguishable by reducing the gap between such domain embeddings, so that it learns domain-invariant representation. Analogous to adversarial learning [21, 18, 41], we alternate the optimization of the domain information extractor and that of the encoder. This approach allows the encoder to gradually become domain-invariant, and the domain information extractor to become sensitive to the remaining domain-specific information of the encoder.
The remaining challenge in implementing TestDG is that only the test domain at hand is accessible at each iteration of CTTA, while domain-invariant learning demands domain embeddings of at least two test domains. To tackle this, TestDG efficiently stores embedding vectors that represent domain embeddings of the previous test domain; we call such embedding vectors domain prototypes and sample them from domain embeddings of the previous domain through submodular optimization [48, 32]. By encouraging that domain embeddings of the current domain are not distinguished from the domain prototypes of the previous domain, TestDG performs domain-invariant learning of the encoder without access to samples of previous test domains.
TestDG was evaluated on the four public benchmarks for CTTA: CIFAR10-to-CIFAR10-C [33, 25], CIFAR100-to-CIFAR100-C [33, 25], ImageNet-to-ImageNet-C [12, 25], and Cityscapes-to-ACDC [11, 56]. It achieved a new state of the art on all the benchmarks and demonstrated greater generalization ability on unseen test domains. The contribution of this paper is three-fold:
  • We propose a new approach to CTTA that learns generalizable features to unseen domains for improving robustness on arbitrary test domains that a model may face in the future.
  • We introduce a novel framework, dubbed TestDG, that learns domain-invariant features during testing without access to previous test domains.
  • TestDG achieved the state of the art on all four public benchmarks and demonstrated great generalization ability on unseen test data.
2 Related work
Test-time adaptation Recent research on TTA aims at addressing the domain shift problem by adapting a pre-trained model online during testing [73, 61, 64, 67, 43, 72]. Initial studies [61, 44] employed self-supervised learning with proxy tasks on unlabeled test data, but they require modifications to the training stage to accommodate the proxy tasks. Meanwhile, Tent [64], a fully test-time adaptation method based on entropy minimization, requires no modifications to the training stage and is applicable to any pre-trained model. Building on this work, several studies [49, 50] have proposed using entropy-based losses in testing environments: EATA [49] utilizes the entropy loss only for reliable samples, while SAR [50] incorporates a sharpness-aware entropy loss to achieve smooth loss landscapes, which are known to improve generalization [16, 66, 7]. Other approaches [30, 59, 44, 15] leveraged lightweight information from source data to facilitate adaptation while mitigating catastrophic forgetting. For example, TTT++ [44] stores mean and covariance matrices, while TTAC [59] stores category-wise features of source data to align feature distributions during testing. Unlike these methods, our method does not require any computation for source data. Instead, it directly extracts domain embeddings from streamed test data, enabling domain-invariant learning during testing.
Continual test-time adaptation In real-world scenarios, testing environments can vary significantly due to multiple factors such as weather, time, and geolocation. Since conventional TTA methods have been studied under the assumption that the test domain is static, they face challenges in continually changing environments. To address this limitation, Wang et al. [67] introduced the first CTTA method, leveraging multiple augmented inputs and a momentum network to generate more reliable pseudo labels. Song et al. [58] and Lee et al. [34] proposed memory-efficient methods that utilized lightweight meta networks and a mixture of experts, respectively. VDP [17] and ViDA [43] utilized visual prompt learning and visual domain adapters, respectively, to disentangle domain-specific and domain-shared knowledge. Continual-MAE [42] improved target domain knowledge extraction through a self-supervised framework. In this context, Yang et al. [71] proposed a versatile network that identifies and refines unreliable pseudo-labels using the source pre-trained model, while Zhu et al. [77] introduced an uncertainty-aware buffering and graph-based constraint. Unlike these methods, TestDG learns domain-invariant features that well generalize to diverse unseen test domains.
Figure 2: Overview of TestDG. When a domain change is detected, TestDG selects domain prototypes that well represent the previous domain embeddings stored in a queue. Each iteration consists of two steps. In Step 1, the domain information extractor $\psi_{\textrm{domain}}$ learns alongside domain amplifier $\psi_{\textrm{amp}}$ to extract domain embeddings from the encoder features using the domain discrimination loss $\mathcal{L}_{\textrm{dis}}$. In Step 2, the encoder $\phi_{\textrm{enc}}$ is trained using the domain-invariant loss $\mathcal{L}_{\textrm{inv}}$ to reduce the gap between domain embeddings and domain prototypes of the previous domain. The domain prototypes are continuously updated to match with the updated model parameters by $\mathcal{L}_{\textrm{update}}$.
Domain generalization Domain generalization (DG) focuses on improving the generalization performance of models on unseen target domains by training them with multiple source domains. Initial work in DG employed domain alignment [41, 57, 28, 39, 68] to learn domain-invariant representations by minimizing the distance of distribution among multiple domains. Recent studies in DG have introduced various strategies, including data augmentation [74, 75, 29], meta learning [3, 14], and representation learning [47, 23]. Inspired by DG, this paper presents a CTTA method that updates an online model through domain-invariant learning to handle continually changing domains.
3 Proposed method
TestDG is an online test-time domain generalization framework for CTTA, designed to learn domain-invariant features from a stream of test-time inputs; its overall pipeline is illustrated in Figure 2. Unlike domain adaptation and generalization, which in general assume that data from at least two domains are available during the entire training process, CTTA allows access to a handful of data from the test domain at hand (i.e., the input data at each test point) for each iteration. This constraint poses a non-trivial challenge for domain-invariant learning.
TestDG addresses this challenge by storing domain embeddings, i.e., embedding vectors containing domain information of the most recent previous domain, in a queue. To prevent excessive memory consumption caused by storing all domain embeddings, we select representative embeddings from the queue of domain embeddings as domain prototypes when the test domain changes.
To achieve domain-invariant learning without corrupting the semantics of features, TestDG operates in two steps as follows. In the first step, the domain information extractor learns to extract domain embeddings, by capturing only domain-specific information from the encoder features. To facilitate this extraction, another extra module called domain amplifier is attached to the encoder for capturing and emphasizing the domain-specific information of the encoder features. We design the domain amplifier as a lightweight adapter structure [10, 27, 52] to minimize additional parameters; details are provided in Sec. B.2 of the supplement. During this step, the domain amplifier is updated alongside the domain information extractor while the encoder remains frozen. In the second step, we perform domain-invariant learning by ensuring the domain embeddings of the current domain become indistinguishable from the domain prototypes of the previous domain. During this step, only the encoder is updated, with the domain information extractor and domain amplifier frozen.
By alternating between the two steps, the encoder gradually learns domain-invariant features across multiple test domains. Meanwhile, the domain information extractor and domain amplifier are updated to reflect the update of the encoder and extract domain embeddings corresponding to the current state of the encoder. In addition, TestDG continuously updates the domain prototypes to match them with the updated model parameters. Since the model used to extract domain prototypes is constantly updated during CTTA, the prototypes must also be updated accordingly.
3.1 Domain change detection
As it is unknown when the test domain changes in CTTA, we detect domain changes based on prediction confidence scores as in [17]. That is, if the difference between confidence scores of two consecutive predictions is greater than a threshold, the test domain is considered to have changed. TestDG is designed to remain effective even with detection errors: false positives are caused by intra-domain shifts and thus let TestDG learn invariant features within a single domain, while false negatives suggest mild distribution shifts that can be handled robustly by the model at hand as-is.
3.2 Domain prototype selection
When a domain change is detected, TestDG retains a small set of $n$ domain prototypes $\mathcal{P}_{\textrm{pre}}=\{p_{i}\}_{i=1}^{n}$ from the previous domain embeddings in the queue $\mathcal{F}_{\textrm{pre}}=\{f_{i}\}_{i=1}^{m}$ to reduce memory footprint substantially. To ensure that these prototypes represent the entire distribution of the previous domain, we select $\mathcal{P}_{\textrm{pre}}$ by minimizing the discrepancy between $\mathcal{F}_{\textrm{pre}}$ and $\mathcal{P}_{\textrm{pre}}$. Following the prototype sampling technique [32], we measure the discrepancy between $\mathcal{F}_{\textrm{pre}}$ and $\mathcal{P}_{\textrm{pre}}$ using the squared maximum mean discrepancy (MMD):
$\textrm{MMD}^{2}(\mathcal{F}_{\textrm{pre}},\mathcal{P}_{\textrm{pre}})$ $\,:=\frac{1}{|\mathcal{F}_{\textrm{pre}}|^{2}}\sum_{f_{i},f_{j}\in\mathcal{F}_{\textrm{pre}}}k(f_{i},f_{j})-\frac{2}{|\mathcal{F}_{\textrm{pre}}||\mathcal{P}_{\textrm{pre}}|}\sum_{f_{i}\in\mathcal{F}_{\textrm{pre}},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(f_{i},p_{j})$
$\,\ \ \ \ \ \ +\frac{1}{|\mathcal{P}_{\textrm{pre}}|^{2}}\sum_{p_{i},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(p_{i},p_{j})\;,$ (1)
where $k(\mathbf{x},\mathbf{x}^{\prime})=\text{exp}(-\gamma||\mathbf{x}-\mathbf{x}^{\prime}||^{2})$ is a RBF kernel that measures the discrepancy between embeddings from the two sets. As the first term in Eq. (1) remains constant with respect to $\mathcal{P}_{\textrm{pre}}$, we define a score function $J(\mathcal{P}_{\textrm{pre}})$ as follows:
$J(\mathcal{P}_{\textrm{pre}})$ $\,:=\text{MMD}^{2}(\mathcal{F}_{\textrm{pre}},\emptyset)-\text{MMD}^{2}(\mathcal{F}_{\textrm{pre}},\mathcal{P}_{\textrm{pre}})$
$\,=\frac{2}{|\mathcal{F}_{\textrm{pre}}||\mathcal{P}_{\textrm{pre}}|}\sum_{f_{i}\in\mathcal{F}_{\textrm{pre}},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(f_{i},p_{j})-\frac{1}{|\mathcal{P}_{\textrm{pre}}|^{2}}\sum_{p_{i},p_{j}\in\mathcal{P}_{\textrm{pre}}}k(p_{i},p_{j})\;,$ (2)
where the constant term $\text{MMD}^{2}(\mathcal{F}_{\textrm{pre}},\emptyset)$ is added to ensure that $J(\emptyset)=0$. The domain prototypes are selected by maximizing the score function:
$\max_{\mathcal{P}_{\textrm{pre}}\subset\mathcal{F}_{\textrm{pre}}:|\mathcal{P}_{\textrm{pre}}|=n}J(\mathcal{P}_{\textrm{pre}})\;.$ (3)
This optimization is intractable, but a near-optimal solution can be achieved through a greedy process since the function $J$ is normalized monotone submodular [48] when using the RBF kernel for $k(\cdot,\cdot)$, as proved in [32]. By maximizing the score function $J(\mathcal{P}_{\textrm{pre}})$, selected prototypes closely approximate the entire distribution of domain embeddings in the queue. At the initial stage where no previous test domain exists, we use domain embeddings of augmented test images as domain prototypes.
3.3 Domain-invariant learning
TestDG updates the model through domain-invariant learning and self-training. For domain-invariant learning, we employ a two-step process. In the first step, the domain information extractor and amplifier are updated to extract domain-specific information, i.e., domain embeddings, from the encoder features. In the second step, the encoder is updated to be domain-invariant by aligning the current domain embeddings with the previous domain prototypes (Sec. 3.2). Alternating these two steps allows the encoder to gradually become domain-invariant, and the domain information extractor and amplifier to become sensitive to the remaining domain-specific information of the encoder. In addition, following previous work [67, 17, 43], we update our model by a self-training loss $\mathcal{L}_{\textrm{self}}$, a cross-entropy loss between our model’s prediction $\hat{y}$ and a pseudo label $\tilde{y}$ generated by its exponential moving average [62]:
$\mathcal{L}_{\textrm{self}}(x)=-\frac{1}{C}\sum_{c=1}^{C}\tilde{y}_{c}\log\hat{y}_{c},$ (4)
where $C$ is the total number of classes. For inference, we use $\hat{y}$ as the classification prediction. Details of the two steps are described below.
Step 1: Extracting domain embeddings. To maintain the semantic information of the encoder features during domain-invariant learning, we first extract their domain-specific information in the form of embedding vectors. To this end, we introduce the domain information extractor that extracts domain embeddings from output features of the encoder, and the domain amplifier that amplifies the domain-specific information of the encoder to facilitate the extractor. Let $\mathcal{F}_{\textrm{cur}}=\{f_{i}\}_{i=1}^{n}$ be the domain embeddings of test input images from the current test domain and $\mathcal{P}_{\textrm{pre}}$ the previous domain’s prototypes. The domain information extractor and amplifier learn a domain embedding space by distinguishing $\mathcal{F}_{\textrm{cur}}$ and $\mathcal{P}_{\textrm{pre}}$ with a domain discrimination loss:
$\mathcal{L}_{\textrm{dis}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})=-\frac{1}{n}\bigg{\{}\displaystyle\sum_{f_{i}\in\mathcal{F}_{\textrm{cur}},}$ $\,\log(D(f_{i}))-\sum_{p_{i}\in\mathcal{P}_{\textrm{pre}}}\log(1-D(p_{i}))\bigg{\}},$ (5)
where $D(\cdot)$ indicates a binary linear classifier that serves as a domain discriminator. Minimizing this loss encourages the domain extractor with the amplifier to learn a space where embeddings of different domains are separated and thus capture domain-specific information [45, 36].
Figure 3: Analysis on the impact of domain-invariant learning in TestDG. (a) 2D visualization of the distribution of domain embeddings from different domains before and after TestDG. (b) The domain gap between different domains measured by Chamfer distance before and after applying TestDG.
Figure 4: Analysis on domain embeddings and prototypes. (a) Analysis of class classification and domain classification accuracy using domain embeddings during CTTA iterations. (b) 2D visualization of domain embeddings and domain prototypes. (c) MMD between the target and original prototypes before the update and that between the target and updated prototypes.
Step 2: Learning domain-invariant features. The second step aims to optimize the encoder to achieve domain invariance across diverse test domains alongside classification (Eq. 4). TestDG accomplishes this by aligning the current domain embeddings and the previous domain’s prototypes. To be specific, we update the encoder and its classification head while freezing the domain information extractor and amplifier. Given current domain embeddings $\mathcal{F}_{\textrm{cur}}$ and previous domain prototypes $\mathcal{P}_{\textrm{pre}}$, the domain-invariant learning loss is given by
$\mathcal{L}_{\textrm{inv}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})=\frac{1}{n}\sum\limits_{(f_{i},p_{i})\in(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})}\left\|f_{i}-p_{i}\right\|_{1}.$ (6)
By minimizing this loss, the encoder’s features are trained to produce domain embeddings that are indistinguishable between different domains, making them domain-generalizable to future domains.
3.4 Domain prototype update
Since the model used to extract the domain prototypes is constantly updated during CTTA, the domain prototypes extracted from the earlier model become outdated. Thus, the domain prototypes should also be updated to reflect the update of the model, approximating prototypes extracted from the updated model. Inspired by Asadi et al. [2], we aim to preserve the set similarity between the previous domain prototypes $\mathcal{P}_{\textrm{pre}}$ and the current domain embeddings $\mathcal{F}_{\textrm{cur}}$ to effectively keep the relationship between them intact. In specific, we measure the Chamfer distance between the two sets $\mathcal{P}_{\textrm{pre}}$ and $\mathcal{F}_{\textrm{cur}}$; a detailed analysis is provided in the supplement. For iteration $t$, we update the domain prototypes $\mathcal{P}_{\textrm{pre}}^{t-1}$ to $\mathcal{P}_{\textrm{pre}}^{t}$ to ensure that the Chamfer distance between previous domain prototypes and current domain embeddings remains consistent before and after the model update. Specifically, the domain prototypes $\mathcal{P}_{\textrm{pre}}^{t}$ are updated by minimizing the prototype updating loss:
$\mathcal{L}_{\textrm{update}}=|d_{\textrm{CD}}(\mathcal{F}_{\textrm{cur}}^{t-1},\mathcal{P}_{\textrm{pre}}^{t-1})-d_{\textrm{CD}}(\mathcal{F}_{\textrm{cur}}^{t},\mathcal{P}_{\textrm{pre}}^{t})|,$ (7)
where the Chamfer distance $d_{\textrm{CD}}$ is computed by
$d_{\textrm{CD}}(\mathcal{F}_{\textrm{cur}},\mathcal{P}_{\textrm{pre}})=\sum_{f_{i}\in\mathcal{F}_{\textrm{cur}}}\min_{p_{j}\in\mathcal{P}_{\textrm{pre}}}\left\|p_{j}-f_{i}\right\|_{2}^{2}+\sum_{p_{j}\in\mathcal{P}_{\textrm{pre}}}\min_{f_{i}\in\mathcal{F}_{\textrm{cur}}}\left\|p_{j}-f_{i}\right\|_{2}^{2}.$ (8)
3.5 Empirical justification
Domain-invariant learning. We analyze the impact of domain-invariant learning both qualitatively and quantitatively. Figure 3(a) provides a $t$-SNE visualization [63] of domain embeddings from different domains (i.e., Gaussian and Shot noise) before and after applying TestDG, demonstrating that TestDG effectively reduces the gap between different domains. Figure 3(b) shows the Chamfer distance between domain embeddings from different domains before and after applying TestDG, indicating that the distance decreases as intended after the domain-invariant learning.
Domain embedding. In Figure 4(a), we empirically verify that domain embeddings capture domain-specific information more prominently than content information. To assess the content information within the domain embeddings, we performed linear probing by training a linear classification layer using the domain embeddings as input while keeping the model frozen. The results show that domain classification using the domain embeddings consistently achieves higher accuracy than class classification, indicating that the domain embeddings predominantly encode domain-specific information with minimal content information, as intended.
Domain prototype selection. We qualitatively investigate how well domain prototypes represent the entire set of domain embeddings. Figure 4(b) provides a $t$-SNE visualization of domain embeddings and selected domain prototypes on the Gaussian noise domain, showing that our domain prototypes are diverse and well represent the overall distribution of the domain embeddings.
Domain prototype update. Figure 4(c) demonstrates the effectiveness of the prototype updating strategy. By forwarding the corresponding images of each prototype through the updated model, we obtain ideal target prototypes in the updated feature space. The MMD value between these target prototypes and our updated prototypes is lower than those between target prototypes and original prototypes, where MMD values are averaged across all CTTA steps. This result validates that our updating method effectively moves the prototypes toward their targeted representations.
4 Experiments
4.1 Experimental setting
Evaluation benchmarks. Our method was evaluated on both classification and semantic segmentation tasks. For classification, we utilized three benchmarks [25]: CIFAR10-C, CIFAR100-C, and ImageNet-C, each consisting of 15 corruption types and five severity levels. We reported the results for the most severe level (level 5). For segmentation, we used Cityscapes [11] as the source and ACDC [56] as a test dataset, including four unseen test domains: Fog, Night, Rain, and Snow.
Evaluation scenario. Following [67], TestDG was evaluated under an online CTTA setting where the prediction is assessed immediately after data are streamed, and the pre-trained model is adapted on the fly. For the segmentation task, test domains are repeated cyclically for three rounds. Additionally, we evaluated the generalization ability to unseen domains on CIFAR10-C using the leave-one-domain-out rule [76, 38, 43], applying CTTA methods to 10 domains of ImageNet-C and testing on the 5 unseen domains without any model adaptation. In line with recent CTTA studies [43, 42, 71], we reported results for all methods using a common ViT backbone. This follows the standardization introduced by ViDA [43], which established the use of ViTs in CTTA and re-implemented previous CNN-based methods within this ViT-based framework. For additional experiments beyond those reported by previous studies, we evaluated methods with publicly available code.
Implementation details. We utilized ViT-base [13] for classification tasks and Segformer-B5 [69] for the segmentation task as the backbone. Input images were resized to $384\times 384$ for CIFAR10-C and CIFAR100-C, $224\times 224$ for ImageNet-C, and to $960\times 540$ for ACDC. The size of the queue storing the domain embeddings and domain prototypes was set to 256 and 40, respectively. The domain information extractor $\psi_{\textrm{domain}}$ and domain discriminator $D$ were both implemented as MLP whose activation functions are ReLU [1] and GELU [26], respectively. The domain amplifier was implemented as an adapter that includes an up-projection layer and a down-projection layer, with an intermediate dimension of 128. More details are given in Sec. B.1 of the supplement.
Table 1: Classification error rates (%) on CIFAR10-to-CIFAR10-C online CTTA task. Gain (%) represents the percentage of improvement in accuracy compared with the source method.
[…]
Table 2: Classification error rates (%) on CIFAR100-to-CIFAR100-C CTTA.
[…]
Source [13] 35.4 /
Pseudo [35] 33.2 +2.2
TENT [65] 32.1 +3.3
CoTTA [67] 34.8 +0.6
VDP [17] 32.0 +3.4
ViDA [43] 27.3 +8.1
Continual-MAE [42] 26.4 +9.0
Versatile [71] 24.0 +11.4
TestDG 23.3 +12.1
Table 3: Classification error rates (%) on ImageNet-to-ImageNet-C CTTA.
[…]
Source [13] 55.8 /
Pseudo [35] 61.2 -5.4
TENT [65] 51.0 +4.8
CoTTA [67] 54.8 +1.0
VDP [17] 50.0 +5.8
ViDA [43] 43.4 +12.4
Continual-MAE [42] 42.5 +13.3
Versatile [71] 42.7 +13.1
TestDG 42.5 +13.3
Table 4: Generalization performance evaluated by error rates (%) on 5 unseen CIFAR10-C domains after applying CTTA over 10 other domains.
[…]
Source [13] 5.3 49.7 23.6 24.7 23.1 25.3 /
ViDA [43] 4.1 36.4 15.6 33.5 17.8 21.5 +3.8
Continual-MAE [42] 5.0 25.7 15.1 22.6 18.7 17.4 +7.9
TestDG 2.5 12.6 12.8 18.7 12.6 11.8 +13.5
Table 5: CTTA results under gradual domain shifts on CIFAR10-to-CIFAR10-C, evaluated by error rates (%).
Source [13] ViDA [43] Continual-MAE [42] TestDG
[…]
Gain / +6.0 +7.2 +10.1
Table 6: Segmentation mIoU results on Cityscape-to-ACDC online CTTA task.
[…]
4.2 Evaluation on classification tasks
To evaluate the performance of TestDG, we conducted extensive experiments on three conventional benchmarks: CIFAR10-C, CIFAR100-C, and ImageNet-C.
Performance on corruption benchmarks. Table 1 and Table 5 present the results on CIFAR10-C and CIFAR100-C. TestDG significantly outperformed all other methods across various corruption types. Specifically, it achieved the lowest mean error rate of 7.7% on CIFAR10-C and 23.3% on CIFAR100-C, showing improvements of 20.5%p and 12.1%p over the source model, respectively. As shown in Table 5, TestDG also achieved the best performance on the ImageNet-C dataset, where performance differences are naturally smaller due to the large number of classes.
Generalization performance on unseen domains. Table 5 summarizes the classification error rates on 5 unseen domains of CIFAR10-C after CTTA over 10 domains. TestDG achieved the lowest mean error rate, highlighting the benefit of domain-invariant learning and demonstrating its generalization ability by maintaining robustness even on unseen domains. This indicates that TestDG not only adapts well to continually changing domains but also generalizes effectively to unseen domains.
Gradually changing results. Table 5 shows results of continual test-time adaptation on a CIFAR10-C gradually changing setup [67], where the corruption types changed gradually over time. TestDG still demonstrated superior performance, notably outperforming existing methods.
4.3 Evaluation on segmentation tasks
Table 6 presents the segmentation results from adapting Cityscapes to ACDC, a real-world adverse condition dataset. Across three rounds, TestDG consistently achieved the highest mIoU among existing CTTA methods. In the first round, our method reached an average mIoU of 62.0%, representing a 5.3%p gain over the source model. In subsequent rounds, TestDG performed strongly on newly encountered domains while preserving robustness on previously encountered domains, effectively handling continually changing environments. These results suggest that TestDG successfully mitigates domain shifts in semantic segmentation tasks, maintaining robustness across diverse and real-world adverse conditions. Further analyses are given in Sec. E.1 and F.3.
Table 7: Ablation study on each component of TestDG.
[…]
21.9
7.7
[…]
28.2
18.0
10.5
8.3
8.0
7.7
Figure 5: Sensitivity to hyperparameters. (a) Threshold for domain change detection. (b) Queue length.
Figure 6: Comparison of computational cost on CIFAR10-C.
4.4 In-depth analysis
Impact of domain amplifier. We investigated the impact of the domain amplifier by evaluating TestDG with and without the domain amplifier by comparing their error rates (Table 4.3). The results show that the domain amplifier plays a crucial role in TestDG, extracting domain-specific information from each layer of the encoder, thereby enabling the encoder to become domain-invariant.
Ablation study of each loss. We conducted an ablation study to evaluate the contributions of each loss component in our method. The results, presented in Table 4.3, demonstrate that each loss contributes to the overall performance improvement. Notably, the domain-invariant learning loss $\mathcal{L}_{\textrm{inv}}$ was crucial for achieving high performance by guiding the model to learn domain-invariant features, which are essential for effective adaptation across different domains. Additionally, the domain discrimination loss $\mathcal{L}_{\textrm{dis}}$, which facilitated the learning of domain embeddings, was also crucial for TestDG. Furthermore, both the score function for domain prototype selection $J(\mathcal{P}_{\textrm{pre}})$ and the prototype updating loss $\mathcal{L}_{\textrm{update}}$ contributed to the performance of TestDG.
Sensitivity to the threshold for domain change detection. To investigate the effect of the threshold used for domain change detection, we conducted experiments varying the threshold values. Figure 5(a) summarizes the results for thresholds ranging from 0.02 to 0.20. The performance remained stable across this range of thresholds, consistently outperforming the source model. These results suggest that TestDG is insensitive to threshold settings.
Sensitivity to the queue length $|\mathcal{F}_{\text{pre}}|$. We evaluated the impact of $|\mathcal{F}_{\text{pre}}|$, which specifies the length of the queue that stores previous domain embeddings. It is designed to retain a sufficient number of embeddings while minimizing the inclusion of excessively outdated information. Figure 5(b) presents the error rates on CIFAR10-C for six different queue lengths. The results indicate that the error rates remain consistent across all tested values of $|\mathcal{F}_{\text{pre}}|$, demonstrating that TestDG is insensitive to the length of the queue. An extended analysis is provided in Sec. C.3.
Comparison of computational cost. We evaluated the computational cost of TestDG in terms of inference latency and memory consumption with the prediction accuracy. As shown in Figure 6, while TENT [64] exhibited the highest computational efficiency, it had the lowest prediction accuracy among the evaluated methods. In contrast, TestDG achieved reduced inference latency and comparable memory consumption compared to other CTTA methods (excluding TENT) while outperforming them in prediction accuracy.
5 Conclusion
We have introduced TestDG, a novel online test-time domain generalization framework for CTTA. Unlike existing CTTA methods that focus only on adaptation to the current test domain and thus often lack generalization to arbitrary future test domains, TestDG aims to learn domain-invariant features during testing through a carefully designed model architecture and test-time training strategy. Through extensive experiments, we have validated that TestDG improves generalization to arbitrary unseen domains, ensures robust adaptation in dynamically changing test environments, and consequently achieves the best on all CTTA benchmarks.
Limitations and future work. TestDG is less robust to initial corruptions due to limited exposure to diverse domain shifts as described in Sec. E.3, and we believe early adaptation strategies could mitigate this issue. Also, TestDG currently uses only the most recent previous domain for domain-invariant learning, and incorporating additional previously encountered domains with minimal memory usage could further improve its performance.